OpenStudy (anonymous):

write the equation of the line in point slope form tht passes through the points (-5, -2) (-4, -5)

sam (.sam.):

First find \[Slope=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

sam (.sam.):

\[Slope=\frac{-2--5}{-5--4}=-3\]

OpenStudy (anonymous):

\[y=mx+b\] \[m = (y_1-y_2)/(x_1-x_2)\] \[m=-3\] \[y=-3x+b\] choose a point \[-5=-3*-4+b\] \[b=-17\] \[y=-3x-17\]

OpenStudy (anonymous):

or alternatively: 

sam (.sam.):

Pick any point, (-5,-2) y-y_{1)=m(x-x_{1}) y+2=-3(x+5) y=-3x-15-2 y=-3x-17

OpenStudy (anonymous):